Vocaloid:

Solving Quadratic Equations Walkthrough for 8bithelix

Vocaloid:

2x^2 + 16x = - 18 first step is to factor the left side such that the x^2 term has a coefficient of 1 2(x^2 + 8x) = -18 divide both sides by 2 to get (x^2 + 8x) = -9 then, to "complete the square" this means we want to find a c term for x^2 + 8x that will make the left hand side a perfect square trinomial to do this, we take the b coefficient (aka the coefficient of the x term, so 8, divide this by 2, square it --> you need to know this step by heart) so that gives us (8/2)^2 = 16 adding this to both sides gives us (x^2 + 8x + 16) = -9 + 16 then simplify to get x^2 + 8x + 16 = 7 then, factor the left side, since it's a perfect square trinomial where (x+a)^2 = x^2 + 2ax + a^2, so: (x+4)^2 = 7 taking the square root of this gives us x + 4 = +/- sqrt(7) (it's a plus minus symbol indicating we are considering the positive and neg values of sqrt(7)) subtracting 4 from each side gives us x = -4 +/- sqrt(7) = your answer x^2 + 8x = 16 = 7

Vocaloid:

(sorry about that last line, idk how that got there)

Vocaloid:

anyway for 6, we just re-arrange the formula in the form ax^2 + bx + c = 0, so it's just adding 1 to each side 2x^2 - 5x + 1 = 0 now, if we compare the quadratic to the general form ax^2 + bx + c, we can see that a = 2, b = -5, and c = 1 (have to be careful about positive/negative signs here) then plug them into the quadratic formula

Vocaloid:



Vocaloid:

this gives us \[\frac{ -(-5 )±\sqrt{(-5)^{2}-4*1*2} }{ 2*2 }\]

Vocaloid:

\[\frac{ 5 )±\sqrt{25-8} }{ 4 }\]

Vocaloid:

then that gives us x = (5±sqrt(17))/4 = your answer

Vocaloid:

lmk if any steps are unclear

8bithelix:

ok